OpenStudy (anonymous):

Slope - Intercept Form ? >_> I chose y=-1/3x+3 & -x+3y=9

OpenStudy (anonymous):

It was in groups . That was group 2 .. I now have to do this : ~ Type the first equation from the group you selected & Identify the slope & The y-intercept of the equation ..

OpenStudy (austinl):

\(\text{Slope Intercept}\rightarrow y=mx+b\)

OpenStudy (anonymous):

I know that -_-

OpenStudy (austinl):

First choice that you made was slope intercept, so I assume that was correct?

OpenStudy (austinl):

The second is in standard form of an equation, \(Ax+By=C\)

OpenStudy (anonymous):

It just says : 1. Choose one of the groups of equations from the table below . Indicate what group you chose . I chose group 2 which is y=1/3x+3 & -x+3y=9

OpenStudy (austinl):

Ok ok, I got it. And now you need to find the two things (slope&y-intercept), right? For slope intercept form it is, \(y=mx+b\\ m=\text{slope}\\ b=\text{y-intercept}\) For the standard form of an line we need to solve it for y to get it into this form.

OpenStudy (anonymous):

First , I wanna know HOW to get the slope & y-intercept ..

OpenStudy (austinl):

For slope-intercept form??

OpenStudy (anonymous):

So is it .. -1/3 for m 3 for b ?

OpenStudy (austinl):

Yep.

OpenStudy (anonymous):

I need the slope .. :I Is the slope -1/3 ?

OpenStudy (austinl):

\(\text{Slope-Intercept}~y=\color{red}{m}x+\color{blue}{b}\) \(\color{red}{m}=\text{slope}\\ \color{blue}{b}=\text{y-intercept}\) \(y=\color{red}{-\dfrac{1}{3}}x+\color{blue}{3}\)

OpenStudy (anonymous):

So that's the slope ? .-.

OpenStudy (austinl):

Yes..............

OpenStudy (anonymous):

Sorry ? I dont understand it .. sheesh .

OpenStudy (anonymous):

y-intercept : (0,3)

OpenStudy (austinl):

Yep.

OpenStudy (anonymous):

How would I graph this using slope - intercept ?

OpenStudy (anonymous):

Go down 3 on the y axis and put a dot ?

OpenStudy (austinl):

Plot a point at (0,3), then go down one, right 3 and plot another point. Then draw a line through the two points.

OpenStudy (anonymous):

I have to change -x+3y=9 into slope - intercept form ..

OpenStudy (austinl):

Correct, just solve it for y.

OpenStudy (anonymous):

Y = 9 though ..

OpenStudy (austinl):

\(-x+3y=9\\ 3y=x+9\\ y=\dfrac{1}{3}x+3\)

OpenStudy (anonymous):

It's the same as the first one . -_-

OpenStudy (austinl):

Nope, it is a positive slope in this one.

OpenStudy (anonymous):

Oh I see ..

OpenStudy (anonymous):

So the slope is y=1/3x+3 & the y-intercept is (0,3) ?

OpenStudy (anonymous):

& To graph that one , You would plot a point of (0,3) , Go up one & To the right 3 ?

OpenStudy (austinl):

slope is \(\dfrac{1}{3}\)

OpenStudy (austinl):

And yep to your last one.

OpenStudy (anonymous):

& Then draw a line through them ?

OpenStudy (austinl):

Yep :)

OpenStudy (anonymous):

Thank you . (:

OpenStudy (austinl):

You are very welcome!